OpenStudy (akohl):

@Abhisar The mass of Saturn is . What is the orbital speed of an ice particle that orbits Saturn at a distance of 102,000 km from the planet's center? Note: The constant of universal gravity (G) equals 6.674 × 10-11 N · m2/kg2. 13.8 km/s 19.3 km/s 1.93 km/s 27.5 km/s Mass should be 5.67*10^26kg

OpenStudy (abhisar):

Equate centripetal force with gravitational force.

OpenStudy (akohl):

There's no mass to get gravitational force

OpenStudy (anonymous):

mass of particle will get cancelled and u have mass of saturn

OpenStudy (akohl):

What

OpenStudy (anonymous):

Gmp/r^2=pv^2/r where p= mass of particle

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\huge\sf \frac{\cancel mV^2}{r}= \frac{G\cancel mM}{r^2}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (akohl):

@abhisar which mass do I use Saturn

OpenStudy (abhisar):

What is this 5.67*10^26kg ?

OpenStudy (akohl):

Saturn's mass

OpenStudy (abhisar):

What do you have to find out ?

OpenStudy (akohl):

The particle's velocity

OpenStudy (abhisar):

Good, then take M=5.67*10^26kg and find V

OpenStudy (akohl):

Is r 102000

OpenStudy (abhisar):

yep, convert it into metres.

OpenStudy (akohl):

I got v^2/r equals 3.64 now what

OpenStudy (akohl):

I got 19268 m/s

OpenStudy (akohl):

Is that 19.3 or 1.93km/s

OpenStudy (abhisar):

\(\huge\sf \frac{\cancel mV^2}{r}= \frac{G\cancel mM}{r^2}\) ---> \(\huge\sf V^2= \frac{GM}{r}\)

OpenStudy (akohl):

I think its 19.3km/s

OpenStudy (akohl):

Is that right?

OpenStudy (abhisar):

I am getting 3.7 * 10^9 m/s

OpenStudy (akohl):

That's not a possible answer though. When I did it gm/r^2 and did v/r without cancelling out the r its 19268m/s but its 3.9*10^9 when you cancel the r

OpenStudy (akohl):

I mean 3.7

OpenStudy (abhisar):

yes u will have to cancel r !

OpenStudy (abhisar):

V^2=GM/r

OpenStudy (akohl):

I don't know then maybe they screwed up the question

OpenStudy (akohl):

Gonna put up another question

OpenStudy (abhisar):

u didn't got ?

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\huge\sf \frac{\cancel mV^2}{\cancel r}= \frac{G\cancel mM}{r\cancel{^2}}\) ---> \(\huge\sf V^2= \frac{GM}{r}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (akohl):

I just put 19.3 bc 3.7*10^9 isn't an option